Exercise 6-4 (Part Level Submission) Comfi Airways, Inc., a small two-plane passenger airline, has asked for your assistance(a1) Your answer is correct. Calculate the break-even point in dollars. Break-even point 33000 $(a2) Calculate the break-even point in number of passenger flights. Break-even point flights
Answer #1

Break even point = Fixed cost/Contribution margin per flight

Contribution margin per flight = 27,360/400 = 68.40

= 18,810/68.40

= 275 flights